7.Wendy's lottery ticket and candy bar budget is ​$ 30 a week.

The price of a lottery ticket is ​$ 2 and the price of a candy bar is ​$5.

Choose the correct statement.

A. When Wendy spends all her budget on candy bars​, she buys 15 candy bars a week.

B. Wendy needs a larger budget.

C. Wendy cannot afford to buy 5 lottery tickets and 4 candy bars a week.

D. Wendy can afford to buy 6 lottery tickets and 5 candy bars a week.

E. When Wendy buys 15 lottery tickets a week she cannot afford to buy a candy bar.
Answer #1

Answer: E. When Wendy buys 15 lottery tickets a week she cannot afford to buy a candy bar.

Option E is correct because if wendy buys 15 lottery tickets @ $2 each, 15*$2 = $30. As the budget is only $30, wendy cannot afford to buy candy bar.

A. When Wendy spends all her budget on candy bars​, she buys 15 candy bars a week.

Wendy cannot afford 15 candy bars @ $5 each, as it is not within the budget.

15 * 5 = $75. (greater than budget)

B. Wendy needs a larger budget. (False)

C. Wendy cannot afford to buy 5 lottery tickets and 4 candy bars a week.

False. Wendy can afford to buy this combination.

(5 * $2) + (4 * $5) = 10 + 20 = $30

D. Wendy can afford to buy 6 lottery tickets and 5 candy bars a week.

Wendy cannot afford this combination as it is not within the budget.

(6 * $2) + (5 * $5) = 12 + 25 = $37 (greater than budget $30)